# IndiBias: A Benchmark Dataset to Measure Social Biases in Language Models for Indian Context

### v1

@@ 5.1 Gender-Religion axis:
Gender-religion intersection bias refers to the specific form of bias that arises from the intersection of an individual’s gender identity and religious affiliation. For our work, we adopt a binary understanding of gender (i.e., male & female) and specifically concentrate on the religious subgroups: Hindu and Muslim. We use first names as the representations of each intersectional identity group such as Hindu-male, Muslim-male, Hindu-female, Muslim-female. We scraped first names from publicly available sources [^5] https://rb.gy/olu2a4 and checked the occurrences of each first name in the pre-training corpus of Muril Khanuja et al. (2021) and IndicBert Doddapaneni et al. (2023) models. We use 14 most frequently occurring names for each intersectional identity group. For the gender-religion intersectional axis, we calculate SEAT scores with Career/Family concepts from Caliskan et al. (2017). The Career/Family word list, as provided by these researchers, serves as our foundation for this analysis. In addition, we extend our investigation by computing SEAT scores using our own Non-violent/Violent concepts (e.g., calm, safe, aggressive, destructive, etc.). For the latter, we have formulated a dedicated word list. Further details, including the list of names used for each intersectional group and the complete Non-violent/Violent word list, are provided in table 9 and 10 respectively in the Appendix. We discuss the bleached sentence patterns to calculate SEAT scores in Appendix A.
@@ 5.2 Gender-Caste axis:
Gender-caste intersection bias refers to the specific form of bias that arises from the intersection of an individual’s gender identity and caste identity. We consider two subgroups corresponding to the caste identity, i.e., lower caste and upper caste. For both lower and higher caste groups, we leverage the terms used by Malik et al. (2022), and the complete word list can be found in Table 11 in the Appendix. We use compound nouns consisting of gendered words and caste terms (e.g., dalit boy, brahmin girl, etc.) as representatives of each intersectional identity group such as lower caste-male, upper caste-male, lower caste-female, upper caste-female.
@@ 5.3 Gender-Age axis:
Gender-age intersection bias refers to the specific form of bias that arises from the intersection of an individual’s gender identity and age group. We consider two subgroups corresponding to age identity, i.e., young people and old people. Here also we use compound nouns consisting of gendered words and age terms (e.g., young boy, old lady, etc.) as representatives of each intersectional identity group such as young-male, old-male, young-female, old-female.
@@
IndiBias dataset is an agglomerate of the Indian CrowS-Pairs, the Indian context-specific attribute tuples, and the bleached sentences for three intersectional axes.
@@ 6.1 Models
Bloom Workshop et al. (2023) Wikipedia YES $7B$
@@
Age 48.34 42.93 50.59 49.88 47.98 45.07 62.03 56.17 60.23 49.37 52.36 48.66 56.87 32.31 45.39 43.77
Disability 63.42 68.57 73.1 55.04 72.94 56.34 74.48 91.17 57.11 61.56 53.26 57.11 50.5 34.83 63.26 60.11
Gender 48.11 53.03 54.86 53.37 49.14 50.06 53.83 55.31 52.91 54.97 50.51 52.8 49.26 38.86 51.89 52.91
Physical-appearance 57.11 51.5 63.19 56.34 73.77 39.38 59.6 58.61 58.97 53.74 60.22 52.53 58.13 49.96 38.39 66.19
Socioeconomic 57.71 66.29 54.1 63.62 60 43.81 60.95 61.52 48.95 53.9 58.29 45.33 58.67 59.24 51.81 60.38
Indian CrowS-Pairs 52.55 (±2.15) 56 (±2.64) 56.1 (±1.17) 55.95 (±2.75) 55.85 (±1.13) 47.1 (±1.65) 58.75 (±1.58) 59.75 (±1.44) 53.75 (±3.94) 54.25 (±2.49) 53.95 (±2.76) 50.55 (±2.29) 53.7 (±2.26) 44.15 (±0.93) 50.4 (±2.85) 55.45 (±1.72)
Table 2: Percentage of sentences in which model favors the corresponding sentence according to stereo-antistereo label in both English and Hindi versions of our dataset. The scores are calculated by averaging the results from five separate runs of the models, each time using a different 80% sample of the dataset. For scores on the Indian CrowS Pair dataset, the standard deviation across these five runs is provided in parentheses. Additionally, scores that are closer to 50 are highlighted in bold for each bias category.
6.2 Evaluation of Indian Crows-Pairs
To measure the likelihood of a sentence, $score(S)$, we calculate the probability of unmodified words conditioned on the modified words $P(U|M,\theta)$ and also the probability of modified words conditioned on the unmodified words $P(M|U,\theta)$. The numbers are reported using $P(U|M,\theta)$ as a measure for $score(S)$, because as mentioned by Nangia et al. (2020), $P(M|U,\theta)$ calculation can be biased by the pre-training data used for training the given model. To calculate the probability $P(U|M,\theta)$, the approximation used by Nangia et al. (2020) is taken into account, i.e the expression used to approximate $P(U|M,\theta)$ is given by
@@ 6.2 Evaluation of Indian Crows-Pairs
where $|C|$ is the number of tokens in the U set.
@@
For models that are primarily encoder-based, $P(U|M,\theta)$ is used as a measure of a score of sentence S, denoted by $score(S)$. For models like mT5, mGPT, and Bloom, which are either decoder-based or seq2seq models, the score of a sentence is calculated based on the probability of the sentence. The probability of a sentence is computed as the sum of the conditional probabilities of each token given all the preceding tokens.
Table 2 represents the results obtained using different models on the IndiBias Dataset. For a given model, we calculate the number of pairs of sentences where $score(S1)$ is greater than $score(S2)$ when the label is stereo (let this count be $n_{1}$), and the number of pairs of sentences where $score(S2)$ is greater than $score(S1)$ when the label is anti-stereo (let this count be $n_{2}$). We then compute the percentage of ($n_{1}+n_{2}$) relative to the total number of sentence pairs. If this percentage is closer to 100, it indicates that the model consistently favors more stereotypical sentences. Conversely, if the value approaches 0, it indicates the model’s preference for anti-stereotypical sentences. An unbiased model would yield a score closer to 50.
For English sentences, Muril achieves a score that is closest to 50 among all other models, whereas for Hindi sentences, mGPT attains a score of 50.4. We also observe that models with scores closer to 50 for English sentences across various bias types do not necessarily correspond to lesser biases in the Hindi language. Also, among all models, mT5 majorly prefers the anti-stereotypical associations across different identities for both English and Hindi.
We also discuss the confidence score in Appendix G and show corresponding plots in figure 5 and 6.
6.3 Intersectional Biases
@@
Test ($\downarrow$) / Model ($\rightarrow$) XLMR IndicBert Muril Bernice mT5 mBART Bloom mGPT Llama-v2 Mistral XLMR IndicBert Muril Bernice mT5 mBART Bloom mGPT Llama-v2 Mistral
@@
Table 3: Intersectional SEAT scores (Effect sizes) for Gender-Religion axis. Large effective scores for each model are in bold. C/F: Career/Family words. N/V: Non-violent/Violent words. The underlined values indicate significance at $p=0.01$.
@@
Test ($\downarrow$) / Model ($\rightarrow$) XLMR IndicBert Muril Bernice mT5 mBART Bloom mGPT Llama-v2 Mistral XLMR IndicBert Muril Bernice mT5 mBART Bloom mGPT Llama-v2 Mistral
@@
Table 3 shows the biases for the gender-religion intersection in English and Hindi for ten multilingual models. Hindi is not there in the pre-training of Llama v2 and Mistral models, hence we do not report scores corresponding to these two models in table 3. We present the results for two types of attributes, namely work (Career/Family) and violence (Non-violence/Violence), as the former is a commonly used stereotype for gender while the latter is for religion Caliskan et al. (2017); Abid et al. (2021). The Career/Family bias between the two genders is higher in India-specific models (IndicBert and Muril) in both English and Hindi, indicating that this particular gender bias may be higher in the Indian context than the Western counterparts. Also, mGPT exhibits significant career/family bias for English sentences. The work bias against the female group is higher in the Muslim religion, while it is slightly lower for Hindu females. As expected, the work bias is very low between the male groups across both religions, while it is interesting to note that it is quite high between Hindu and Muslim females in the Hindi models. The violence bias is usually against the Muslim group in all the models across languages. However, Hindi models show higher violence bias against Muslim group.
@@
In the case of gender-caste intersectional bias (Table 4), most of the English models are usually biased toward the female group in terms of their pleasantness, with the exception that Bernice, IndicBert, and Muril illustrate bias toward the upper caste groups when comparing genders across the castes. It is the opposite in Hindi – the models are biased toward the male groups for pleasantness. It is interesting to note that when comparing castes keeping the gender constant, almost all the models for English and Hindi indicate more pleasantness toward the upper caste groups, whereas mBART is biased toward the lower castes, in both languages. The primary factor contributing to this phenomenon can be attributed to the composition of pre-training data used in various models. As evident from the data presented in table 1, the Bernice model’s pre-training involves social media content, where posts on Indian social media platforms exhibit a notable trend of positivity towards individuals belonging to the upper caste, as opposed to those from the lower castes. Similarly, both IndicBERT and Muril models draw their pre-training data from Indian news articles and Indian wiki pages, which consistently display a higher prevalence of positive sentiments directed towards upper-caste individuals compared to those from lower castes.
@@ 7 Conclusion and Future Work
Through IndiBias, we aim to facilitate advancements in the understanding of social biases in LMs, with a specific focus on Indian languages and cultural contexts. We follow a translate-filter-modify approach to create an Indian version of the CrowS-Pairs dataset in English and Hindi languages. Additionally, we have released an extensive set of identity-attribute tuples, encompassing seven different demographics such as gender, religion, caste, age, region, physical appearance, and occupation, to capture positive and negative stereotypes prevalent in Indian society. In addition, our analysis using SEAT revealed the existence of intersectional biases in the Indian context. This finding highlights the significance of considering the compounded effects of multiple dimensions in LM biases. Additionally, we aim to augment the dataset by incorporating sexual orientation instances into the Indian CrowS-Pairs. Also, we intend to expand the tuple dataset by generating template-based sentences that can be incorporated into the Indian CrowS-Pairs.
@@
Appendix A Embedding Association Test
@@
Appendix B Annotator Demographics
@@
Table 5: Prompts used to generate Indian stereotypes. This is referred in the Section 4.2: Bias tuple creation (4.2) in the main paper.
Appendix C Prompts Used

### v2

@@ 5.1 Gender-Religion axis:
Gender-religion intersection bias refers to the specific form of bias that arises from the intersection of an individual’s gender identity and religious affiliation. For our work, we adopt a binary understanding of gender (i.e., male & female) and specifically concentrate on the religious subgroups: Hindu and Muslim. We use first names as the representations of each intersectional identity group such as Hindu-male, Muslim-male, Hindu-female, Muslim-female. We scraped first names from publicly available sources [^10] https://rb.gy/olu2a4 and checked the occurrences of each first name in the pre-training corpus of Muril Khanuja et al. (2021) and IndicBert Doddapaneni et al. (2023) models. We use 14 most frequently occurring names for each intersectional identity group. For the gender-religion intersectional axis, we calculate SEAT scores with Career/Family concepts from Caliskan et al. (2017). The Career/Family word list, as provided by these researchers, serves as our foundation for this analysis. In addition, we extend our investigation by computing SEAT scores using our own Non-violent/Violent concepts (e.g., calm, safe, aggressive, destructive, etc.). For the latter, we have formulated a dedicated word list. Further details, including the list of names used for each intersectional group and the complete Non-violent/Violent word list, are provided in table 10 and 11 respectively in the Appendix. We discuss the bleached sentence patterns to calculate SEAT scores in Appendix B.
@@ 5.2 Gender-Caste axis:
Gender-caste intersection bias refers to the bias that arises from the intersection of an individual’s gender identity and caste identity. We consider two subgroups corresponding to the caste identity, i.e., lower caste and upper caste. For both lower and higher caste groups, we leverage the terms used by Malik et al. (2022), and the complete word list can be found in Table 12 in the Appendix. We use compound nouns consisting of gendered words and caste terms (e.g., dalit boy, brahmin girl, etc.) as representatives of each intersectional identity group such as lower caste-male, upper caste-male, lower caste-female, upper caste-female.
@@ 5.3 Gender-Age axis:
Gender-age intersection bias refers to the bias that arises from the intersection of an individual’s gender identity and age group. We consider two subgroups corresponding to age identity, i.e., young people and old people. Here also we use compound nouns consisting of gendered words and age terms (e.g., young boy, old lady, etc.) as representatives of each intersectional identity group such as young-male, old-male, young-female, old-female.
@@
IndiBias dataset is an agglomerate of the Indian CrowS-Pairs (ICS), the Indian context-specific attribute tuples, and the bleached sentences for three intersectional axes.
@@ 6 Experiments and Results
Bloom Workshop and Team (2023) Wikipedia YES $7B$
@@
6.1 Evaluation of Indian Crows-Pairs
To measure the likelihood of a sentence, $score(S)$, we calculate the probability of unmodified words conditioned on the modified words $P(U|M,\theta)$ and also the probability of modified words conditioned on the unmodified words $P(M|U,\theta)$. The numbers are reported using $P(U|M,\theta)$ as a measure for $score(S)$, because as mentioned by Nangia et al. (2020), $P(M|U,\theta)$ calculation can be biased by the pre-training data used for training the given model.
@@ 6.1 Evaluation of Indian Crows-Pairs
Age 49.69 43.85 51.62 39.93 49.25 40.26 54.25 50.67 65.3 53.22 58.32 54.26 56.17 44.18 58.01 54.62
Disability 75.69 91.91 83.98 58.49 72.01 33.48 75.79 88.13 74.62 62.34 57.63 62.94 53.58 47.89 61.67 85.05
Gender 52.55 53.88 56.84 58.67 52.04 45.82 55.61 58.78 54.29 54.08 51.53 52.35 52.76 40.31 53.47 51.53
Physical-appearance 51.82 50.39 67.65 67.88 70.06 29.46 64.4 66.81 55.27 45.19 63 47.95 51.43 48.7 50.29 60.85
Socioeconomic 61.12 63.55 51.78 45.98 57.76 49.16 70.47 73.64 49.16 52.52 54.77 48.79 56.26 56.82 53.08 63.93
Religion 61.25 63.47 48.95 61.74 59.51 52.22 59.27 59.28 59.52 46.63 49.18 59.6 62.01 55.19 53.7 51.98
Caste 44.07 34.14 42.48 49.88 45.76 56.35 51.05 53.97 50.76 54.8 57.6 61 52.35 37.08 56.45 55.98
ICS (mean) 55.32 55.64 54.86 54.54 55.57 45.96 59.93 62.21 55.89 52.36 54.21 53.79 55.04 46.68 54.32 56.79
ICS (std-dev) (±2.26) (±1.85) (±2.59) (±3.37) (±1.01) (±0.87) (±1.77) (±2.29) (±1.5) (±1.68) (±1.78) (±2.98) (±2.28) (±1.97) (±2.26) (±2.3)
Table 2: Bias Percentage of different models on ICS dataset (as described in section 6.1.1). The scores are calculated by averaging the results from five separate runs of the models, each time using a different 80% sample of the dataset. The standard deviation across these five runs is provided in parentheses. Scores closer to 50 represent that the model is least biased, and such scores are highlighted in bold for each bias category.
To calculate the probability $P(U|M,\theta)$, the approximation used by Nangia et al. (2020) is taken into account, i.e the expression used to approximate $P(U|M,\theta)$ is given by
@@
where $|C|$ is the number of tokens in the U set. $score(S)$ is then calculated by normalizing the probability $P(U|M,\theta)$ relative to the number of tokens $|C|$.
@@
For models that are primarily encoder-based, $P(U|M,\theta)$ is used as a measure of a score of sentence S. For models like mT5, mGPT, and Bloom, which are either decoder-based or seq2seq models, the score of a sentence, $score(S)$ is calculated based on the normalized probability of the sentence. The normalized probability of a sentence is calculated by dividing the sum of the conditional log probabilities of each token, conditioned upon all preceding tokens, by the total number of tokens.
6.1.1 Bias Percentage Calculation
Table 2 represents the results obtained using different models on the IndiBias Dataset. For a given model, we calculate the number of pairs of sentences where $score(S1)$ is greater than $score(S2)$ when the label is stereo (let this count be $n_{1}$), and the number of pairs of sentences where $score(S2)$ is greater than $score(S1)$ when the label is anti-stereo (let this count be $n_{2}$), and we term this as Bias Percentage of the model. We then compute the percentage of ($n_{1}+n_{2}$) relative to the total number of sentence pairs. If this percentage is closer to 100, it indicates that the model consistently favors more stereotypical sentences. Conversely, if the value approaches 0, it indicates the model’s preference for anti-stereotypical sentences. An unbiased model would yield a score closer to 50.
For English sentences, Bernice, IndicBERT, and mT5 achieve scores that are closer to 50 compared to other models. In contrast, for Hindi sentences, XLMR attains a score of 52.36. This observation suggests that models with scores closer to 50 for English sentences across various bias types do not necessarily translate to reduced biases in Hindi. Notably, mT5 predominantly favors anti-stereotypical associations for both English and Hindi. We can also observe that models generally exhibit more bias in English compared to Hindi on the overall ICS dataset. This can be attributed to the difference in the language-specific pre-training corpus for different models, particularly in capturing stereotypes within the Indian context.
Within the gender category, which constitutes the highest percentage of sentences in our dataset, mBART shows the least bias in English, whereas Bloom exhibits the least bias in Hindi. For the religion bias, generally, the models are more biased in English than Hindi, potentially because the English language pre-training corpus captures the concept of religious bias more globally rather than being limited to the Indian context.
We also discuss the difference of scores assigned by models to the pairs of sentences in Appendix C and show corresponding distribution plots in figure 4 and 5.
@@ 6.1.1 Bias Percentage Calculation
Test ($\downarrow$) / Model ($\rightarrow$) XLMR IndicBert Muril Bernice mT5 mBART Bloom mGPT Llama-v2 Mistral XLMR IndicBert Muril Bernice mT5 mBART Bloom mGPT
@@
Table 3: Intersectional SEAT scores (Effect sizes) for Gender-Religion axis. Large effective scores for each model are in bold. C/F: Career/Family words. N/V: Non-violent/Violent words. The underlined values indicate significance at $p=0.01$. NOTE: Hindi language data are not there in the pre-training corpus of Llama-v2 and Mistral.
@@
Test ($\downarrow$) / Model ($\rightarrow$) XLMR IndicBert Muril Bernice mT5 mBART Bloom mGPT Llama-v2 Mistral XLMR IndicBert Muril Bernice mT5 mBART Bloom mGPT
@@
6.2 Evaluation of Intersectional Biases
Table 3 shows the biases for the gender-religion intersection in English and Hindi for ten multilingual models. Hindi is not there in the pre-training of Llama v2 and Mistral models, hence we do not report scores corresponding to these two models in table 3. We present the results for two types of attributes, namely work (Career/Family) and violence (Non-violence/Violence), as the former is a commonly used stereotype for gender while the latter is for religion Caliskan et al. (2017); Abid et al. (2021). The Career/Family bias between the two genders is higher in India-specific models (IndicBert and Muril) in both English and Hindi, indicating that this particular gender bias may be higher in the Indian context than the Western counterparts. Also, mGPT exhibits significant career/family bias for English sentences. The work bias against the female group is higher in the Muslim religion, while it is slightly lower for Hindu females. As expected, the work bias is very low between the male groups across both religions, while it is interesting to note that it is quite high between Hindu and Muslim females in the Hindi models. The violence bias is usually against the Muslim group in all the models across languages. However, Hindi models show higher violence bias against Muslim groups.
@@ 6.2 Evaluation of Intersectional Biases
In the case of gender-caste intersectional bias (Table 4), most of the English models are usually biased toward the female group in terms of their pleasantness, with the exception that Bernice, IndicBert, and Muril illustrate bias toward the upper caste groups when comparing genders across the castes. It is the opposite in Hindi – the models are biased toward the male groups for pleasantness. It is interesting to note that when comparing castes keeping the gender constant, almost all the models for English and Hindi indicate more pleasantness toward the upper caste groups, whereas mBART is biased toward the lower castes, in both languages. The primary factor contributing to this phenomenon can be attributed to the composition of pre-training data used in various models. As evident from the data presented in table 1, the Bernice model’s pre-training involves social media content, where posts on Indian social media platforms exhibit a notable trend of positivity towards individuals belonging to the upper caste Kain et al. (2021), as opposed to those from the lower castes. Similarly, both IndicBERT and Muril models draw their pre-training data from Indian news articles and Indian wiki pages, which consistently display a higher prevalence of positive sentiments [^11] https://shorturl.at/mKMU3 directed towards upper-caste individuals compared to those from lower castes Fonseca et al. (2019); Kureel (2021).
@@ 7 Conclusion and Future Work
Through IndiBias, we aim to facilitate advancements in the understanding of social biases in LLMs, with a specific focus on Indian languages and cultural contexts. We have released an extensive set of identity-attribute tuples encompassing seven different demographics such as gender, religion, caste, age, region, physical appearance, and occupation, to capture positive and negative stereotypes prevalent in Indian society. We follow a translate-filter-modify approach to create an Indian version of the CrowS-Pairs dataset in English and Hindi languages. We then augment this dataset using manually annotated sentence pairs using the tuple dataset. We conducted a comprehensive bias analysis of different LMs using this dataset. In addition, our analysis using SEAT revealed the existence of intersectional biases in the Indian context. This finding highlights the significance of considering the compounded effects of multiple dimensions in LM biases. Additionally, we aim to augment the dataset by incorporating sexual orientation instances into the Indian CrowS-Pairs. Also, we intend to expand such dataset to multiple Indian languages.
Acknowledgements
We would like to thank all our annotators for helping us to create this benchmark dataset. We also thank our anonymous reviewers as well as the ARR, NAACL action editors. Their insightful comments helped us improve the current version of the paper.
@@ Acknowledgements
Appendix A Experimental Setup
Appendix B Embedding Association Test
@@
Appendix C Other Experiments
$DS=\begin{cases}\text{score($S_{1}$)}-\text{score($S_{2}$)},&\text{if stereo}.\\ \text{score($S_{2}$)}-\text{score($S_{1}$)},&\text{if antistereo}.\end{cases}$ (4)
The difference of scores ($DS$) for a given pair of sentences is calculated as $score(S_{1})-score(S_{2})$ for stereo-labeled sentence pairs and $score(S_{2})-score(S_{1})$ for antistereo-labeled sentence pairs. A distribution centered closely around zero suggests that the model exhibits minimal variance among the calculated difference of scores. It represents that, on average, the model does not disproportionately favor one sentence over the other in most instances. Figure 4 and figure 5 show the KDE plot distribution obtained for the difference of scores using English and Hindi sentence pairs of the ICS dataset. For Hindi, mGPT exhibits the highest concentration of difference scores around zero, whereas for English, Bernice demonstrates the highest density of difference scores for the same region. The models exhibit a broader range of differences of scores for English sentence pairs compared to Hindi sentence pairs. Also, we notice that among all models, the distribution for mT5 is skewed towards the negative side for both Hindi and English, thus confirming the bias percentage score of the model (as defined in section 6.1.1) being less than 50 as presented in Table 2.
Figure 4: KDE-Plot of difference of scores ($DS$) for English Sentence pairs in ICS dataset.
Figure 5: KDE-Plot of difference of scores ($DS$) for Hindi Sentence pairs in ICS dataset.
Appendix D Annotator Demographics
@@
Our task is to create a Benchmark dataset for examining the presence of bias and stereotypes in Language models.
For the same we need to generate sentences, which will be validated by humans and used in the dataset.
Some generated sentences can be biased and offensive, but it is the required in order to identify and mitigate biases in LLMs.
I have tuples of the type (identity, attribute), where identity describes
the religion of a person, and attribute is a stereotypically associated attribute with people of that religion.
Now I want you to generate 5 naturally occurring sentences about the tuple (Muslim, extremist).
A naturally occurring sentence is one that occurs in novels, human conversations, movie dialogues, etc. chat/gpt-3.5-turbo
I have tuples of the type (identity, attribute), where identity describes the religion of a person,
and attribute is a stereotypically associated attribute with people of that religion.
Now I want you to generate 5 naturally occurring sentences about the tuple (Muslim, narrow-minded).
A naturally occurring sentence is one that occurs in novels, human conversations, movie dialogues, etc.
This task is to create a Benchmark dataset for examining the presence of bias and stereotypes in Language models,
so some generated sentences can be biased and offensive, but it is the required in order to identify and mitigate biases in LLMs. chat/gpt-3.5-turbo
Table 5: Prompts used to generate Indian stereotypes and to generate sentences from stereotypical tuples. This is referred in the Section 4.2: Bias tuple creation (4.1) and Section 4.2.2 Using IndiBias tuples 4.2.2 respectively in the main paper.
Appendix E Indian Crows-Pairs Statistics
The detailed statistics of the ICS dataset are provided in table 6. The average length of both English and Hindi sentences is more in the pairs annotated using the Tuples than in the pairs created by inheriting Nangia et al. (2020).
Category Overall Inhereted Tuples
Avg. Word Len English 12.77 12.01 15.27
Avg. Word Len Hindi 15.54 14.48 19.08
Stereo/Antistereo Dist (in percentage)
Stereo 81.84 88.00 61.35
Antistereo 18.16 12.00 38.65
Bias Type Dist (in percentage)
Gender 35.03 45.57 -
Socioeconomic 19.14 24.90 -
Religion 14.46 - 62.57
Age 11.06 14.39 -
Caste 8.65 - 37.42
Physical-appearance 7.37 9.59 -
Disability 4.25 5.53 -
Table 6: Overview of IndiBias Dataset Statistics: This table provides details of dataset’s composition, featuring average word lengths in English and Hindi, distribution of stereo and antistereo content, and a breakdown of bias types. The "Overall" column provides comprehensive statistics, while the "Inherited" and "Tuples" columns focus on specific subsets of ICS dataset as detailed in sections 4.2.1 and 4.2.2, respectively.
Appendix F Prompts Used